Randy is four times as old as Marie. In 12 yr, Marie's age will be half of Randy's. How old are they now?
Answer by jorel1380(3719)

R=4M
R+12=2(M+12)
R+12=2M+24
4M-2M=24-12
2M=12
M=6
R=24.